# Learning to Adapt: Bio-Inspired Gait Strategies for Versatile Quadruped Locomotion

### v1

@@ Abstract
Deep reinforcement learning (DRL) has revolutionised quadruped robot locomotion, but existing control frameworks struggle to generalise beyond their training-induced observational scope, resulting in limited adaptability. In contrast, animals achieve exceptional adaptability through gait transition strategies, diverse gait utilisation, and seamless adjustment to immediate environmental demands. Inspired by these capabilities, we present a novel DRL framework that incorporates key attributes of animal locomotion: gait transition strategies, pseudo gait procedural memory, and adaptive motion adjustments. This approach enables our framework to achieve unparalleled adaptability, demonstrated through blind zero-shot deployment on complex terrains and recovery from critically unstable states. Our findings offer valuable insights into the biomechanics of animal locomotion, paving the way for robust, adaptable robotic systems.
@@
Figure 1: Instilling the core animal locomotion proficiency attributes within a DRL locomotion framework. From taking an abstracted view of animal locomotion, the attributes of proficient locomotion are gait transition strategies, gait procedural memory, and adaptive motion adjustment. Taking on a similar structure, within the DRL locomotion framework these attributes are realised by the gait selection policy, $\pi_{G}$, bio-inspired gait scheduler, and locomotion policy, $\pi_{L}$. $\pi_{G}$ has been trained to minimse the animal gait transition metrics applied to the quadruped robot based on the current robot state, $\bm{s}$, and relevant bio-inspired gait scheduler output, $\bm{\beta}_{G}$, to select the optimal gait, $\Gamma^{*}$. The bio-inspired gait scheduler then generates the gait references informed by $\bm{s}$ from encoded high-level gait parameters. The gait references, $\bm{\beta}_{L}$, are then passed to $\pi_{L}$ to inform of any adjustments to the nominal gait motions.
Originally inspired by the remarkable adaptability of quadruped mammal locomotion—an ability shaped by innate and environmentally induced factors [1, 2, 3]—the field of quadruped robotics has made significant efforts to develop equally proficient locomotion frameworks. Presently, the most advanced systems rely on end-to-end deep reinforcement learning (DRL), which involves training a multilayer perceptron (MLP) [4] capable of navigating diverse environments. These frameworks demonstrate robustness in traversing real-world [5] and urban terrains [6], resisting perturbations [7], jumping between platforms [8], overcoming deformable surfaces [9], and recovering from falls [10]. Despite these achievements, their adaptability remains constrained, as most systems are limited to deploying a single targeted gait or locomotion strategy.
In contrast, biomechanics research has shown that no single gait is universally effective across all scenarios [11, 12, 13]. Animals adapt their locomotion by employing nominal gaits such as ambling, trotting, and running [14], while switching to specialised gaits like hopping, pronking, and bounding for off-nominal tasks such as predator evasion or obstacle navigation [15]. Current DRL frameworks fall short of replicating this level of versatility. To address this limitation, some approaches have focused on training DRL policies to learn multiple gaits by providing reference motions during training [16, 17, 18, 19], or by learning from policies that specialise in specific gaits [20]. However, these methods remain insufficient when compared to the extensive capabilities observed in animal locomotion, which include:
@@
While bio-inspired methods that leverage central pattern generators (CPGs) [21] have realised spontaneous gait transitions [22] and mimic certain animal behaviours [23], their performance in real-world applications is often limited. Typically, such experiments are constrained to controlled environments [24, 23] or, when conducted outdoors, are restricted by low velocities and simple tasks [18]. These limitations suggest that instead of attempting to replicate animal locomotion mechanisms precisely, state-of-the-art DRL frameworks should be augmented with high-level attributes derived from animal locomotion to instil the proficiency observed in nature.
@@
Despite these insights, there has been no prior attempt to simultaneously integrate all these attributes—gait transition strategies, procedural memory, and adaptive motion adjustments—within a DRL locomotion framework. This leads to the following research questions:
@@
To address these questions, we propose a novel locomotion framework (see Fig. 1) designed to incorporate these key animal locomotion attributes. The framework demonstrates exceptional adaptability through zero-shot deployment in complex, real-world environments, relying solely on intra-perceptive sensors.
@@ I Results
We evaluate the proposed framework through a set of studies, demonstrating it outperforms others by exhibiting quadruped animal locomotion strategies, and validating this gained proficiency on real-world terrain, as presented in Fig. 2 and Supplementary Video 1 (Appendix B).
@@ I-A Achieving Adaptive Motion Adjustment with a Diverse Set of Gaits
Figure 3: Locomotion comparison study experiments (a) each policy follows a set of command velocities in $x$ and yaw between 00 to $1$ m/s and $-1$ to $1$ rad/s respectively. During each velocity pair, the commanded gait is cycled through all gaits, switching every $1$ s. This repeated $5$ times over flat, $h_{\text{terr}}^{0}$, to very rough terrain, $h_{\text{terr}}^{3}$ (shown in Fig. 2) with the average performance being plotted. A number rather than a magnitude indicates the count of experiments that the policy failed. (b) each policy follows a sinusoidal type trajectory in $v_{x}$ and $\omega_{z}$ while switching gaits every $2$ s, which is repeated $5$ times over $h_{\text{terr}}^{0}$ and $h_{\text{terr}}^{3}$ terrain, with blue highlighted areas indicating transition phases and black triangles representing points where the robot failed. (c) policies follow a command of just $v_{x}=0.5$ m/s while using a trot gait while encountering rectangular steps with a height of $0.05$ m.
@@
To evaluate our method of instilling adaptive motion adjustment and procedural memory for diverse gait deployment, a comparison study is completed between our bio-inspired locomotion policy $\pi_{L}^{\text{bio}}$, a standard multi-gait locomotion policy with no pseudo procedural memory $\pi_{L}^{\text{no}\bm{\beta}_{L}}$, and a policy trained which also uses $\bm{\beta}_{L}$ within $\bm{o}_{L}$ but implements the standard approach of extracting the observations from the simulator, $\pi_{L}^{\text{noSE}}$, for which a video summary can be found in Supplementary Video 2 (Appendix C). From the results of this study, shown in Fig. 3, the proficiency of $\pi_{L}^{\text{bio}}$ over $\pi_{L}^{\text{no}\bm{\beta}_{L}}$ and $\pi_{L}^{\text{noSE}}$ in terms of velocity tracking error $v_{B}^{\text{err}}$, contact schedule tracking error $c_{\text{avg}}^{\text{err}}$, and base stability (magnitude of undesirable base angular velocities) $\omega_{B}^{\text{err}}$, is stark.
As illustrated by Fig. 3(a), on flat terrain $\pi_{L}^{\text{bio}}$ exhibits lower $v_{B}^{\text{err}}$, $c_{\text{avg}}^{\text{err}}$ and $\omega_{B}^{\text{err}}$ compared to $\pi_{L}^{\text{no}\bm{\beta}_{L}}$ and $\pi_{L}^{\text{noSE}}$, on average by $15\%$, $21\%$ and $10\%$ respectively (not accounting for failure cases), with this error increasing at higher velocity command magnitudes. However, this difference in performance is only exacerbated when rough terrain is introduced, in which both $\pi_{L}^{\text{no}\bm{\beta}_{L}}$ and $\pi_{L}^{\text{noSE}}$ fail repeatedly, particularly at higher velocities and rougher terrain as indicated within Fig. 3(a) by the failure occurrence counts within the heatmap, whereas $\pi_{L}^{\text{bio}}$ completes every experiment without fail despite only having experienced flat terrain during training. In turn, this demonstrates its impressive adaptability to new environments (see Section III-C for justification of omitting rough terrain from $\pi_{L}$ training).
@@
This incompetence of $\pi_{L}^{\text{no}\bm{\beta}_{L}}$ and $\pi_{L}^{\text{noSE}}$ is caused by a lack of adaptive swing foot motion adjustments captured within $\bm{\beta}_{L}$ based on $\bm{s}$ and the accumulation of error within the SE respectively. Both of these factors are substantially affected by the instabilities rough terrain inflicts upon the robot. Considering that the nominal swing foot peak height is defined as $25\%$ of the nominal base height and for $h_{\text{terr}}^{2}$ and $h_{\text{terr}}^{3}$ (as defined in Fig. 2) the peak terrain height is $44\%$ and $67\%$ of the base height, having no data or strategy to account for this harsh terrain results in the rapid deterioration of proficiency.
@@
For $\pi_{L}^{\text{no}\bm{\beta}{L}}$, Fig. 3(b) shows large spikes in $c_{\text{avg}}^{\text{err}}$ and $\omega_{B}^{\text{err}}$ when encountering $h_{\text{terr}}^{3}$, highlighting its inability to adapt swing foot trajectories and overcome an unrefined solution space. Fig. 3(c) further shows sustained instabilities in base height, roll, and pitch after contact with steps at $17\%$ of the nominal base height. For $\pi_{L}^{\text{noSE}}$, Figs. 3(b) and 3(c) show that poor reference tracking and stability worsen with velocity command magnitude and time, as it lacks strategies to mitigate error build-up in the SE. With $\pi_{L}^{\text{bio}}$ not experiencing any of these limitations, this explicitly demonstrates the effectiveness of implementing $\bm{\beta}_{L}$ and $\bm{s}$ within $\bm{o}_{L}$; $\pi_{L}^{\text{bio}}$ can deploy a diverse set of gaits while exhibiting highly adaptive behaviour over previously unobserved terrain, in turn demonstrating successful instillation of adaptive motion adjustment and gait procedural memory.
@@ I-B Applying Biomechanics Metrics For Optimal Gait Selection
Directly applying biomechanics metrics to instil animal gait transition strategies is unsuitable due to differences between animals and robots and $\pi_{G}$ training requirements. Therefore, we instead apply cost of transport, CoT, torque saturation, $\tau_{\%}$, external work, $W_{\text{ext}}$, and foot contact reference tracking error, $c_{\text{avg}}^{\text{err}}$, for the minimisation of energy consumption, musculoskeletal forces (generically referred to as actuator-structural forces), mechanical work, and instability respectively. Full details and justification of these adapted metrics are detailed in Section III-E.
@@
In accordance with Section III-F, these metrics are unified within the training of the gait selection policy $\pi_{G}^{\text{uni}}$ to instill the strategies animals use for optimal gait selection for exemplary adaptability. To investigate whether $\pi_{G}^{\text{uni}}$ effectively minimises these metrics through gait selection, the results of competing the highly demanding velocity command trajectory presented in Fig. 4 and Supplementary Video 3 (Appendix D) for $\pi_{G}^{\text{uni}}$ paired with $\pi_{L}^{\text{bio}}$ are collected, along with that for all individual gaits deployable by $\pi_{L}^{\text{bio}}$, for flat terrain and terrain $h_{\text{terr}}^{2}$.
@@
From the gait selection results in the bottom time series subplot presented in Fig. 4, it is clear that $\pi_{G}^{\text{uni}}$ almost exclusively utilises trotting at low speeds and running at high speeds. When accelerating from low to high speeds, $\pi_{G}^{\text{uni}}$ oscillates between trotting and running, with an increasing bias towards running, to increase the stride frequency as visualised by the foot contact data in Fig. 4. Consequently, $\pi_{G}^{\text{uni}}$ not only reliably tracks the optimal gait but outperforms individual gaits in metrics and velocity tracking by oscillating between trotting and running during acceleration. This behaviour, although never targeted during training, is reflected in animal locomotion strategies, where gait stride frequency increases with speed, and transitional phases blend gaits to minimize energy use [29].
However, introducing rough terrain and high acceleration causes $\pi_{G}^{\text{uni}}$ to utilise other gaits as auxiliary tools to overcome instability. Hence a gait classification arises: trot and run gaits are nominal performance gaits at slower and faster speeds respectively, while bound, pronk, limp, amble and hop gaits are auxiliary gaits for off-nominal scenarios such as stability recovery.
@@
When inspecting the radar charts in Fig. 4, which depict $\pi_{G}^{\text{uni}}$ and each gaits relative performance in terms of the metrics, the origin of this emerged gait selection strategy becomes clear. On flat terrain, trot and run gaits achieve a good performance across the metrics with the gaits either exhibiting relatively worse performance. However, when it comes to overcoming rough terrain bound, hop and limp gaits all gain relative performance in $\tau_{\%}$, $W_{\text{ext}}$ and $c_{\text{avg}}^{\text{err}}$, while trot and run gaits exhibit a reduced dominance in relative proficiency. In addition to this observation providing an insight as to why $\pi_{G}^{\text{uni}}$ chooses to utilise these auxiliary gaits it also provides evidence to suggest that $\tau_{\%}$, $W_{\text{ext}}$ and $c_{\text{avg}}^{\text{err}}$ can effectively characterise stability. This observation is further investigated and discussed in the following sections. Overall, $\pi_{G}^{\text{uni}}$ outperforms all individual gaits across all metrics, with the exception of CoT for trot and run gaits where the difference is negligible, demonstrating the successful minimisation of the metrics and successful instillation of gait procedural memory of how to utilise each gait given the robot’s state and task.

### v2

# Learning to Adapt through Bio-Inspired Gait Strategies for Versatile Quadruped Locomotion

@@ Abstract
Deep reinforcement learning (DRL) has revolutionised quadruped robot locomotion, but existing control frameworks struggle to generalise beyond their training-induced observational scope, resulting in limited adaptability and gait proficiency. In contrast, animals achieve exceptional adaptability through gait transition strategies, diverse gait utilisation, and seamless adjustment to immediate environmental demands. Inspired by these capabilities, we present a novel DRL framework that incorporates key attributes of animal locomotion: gait transition strategies, pseudo gait procedural memory, and adaptive motion adjustments. This approach enables our framework to achieve unparalleled adaptability, demonstrated through blind zero-shot deployment on complex terrains and recovery from critically unstable states. Our findings offer valuable insights into the biomechanics of animal locomotion, paving the way for robust, adaptable robotic systems.
@@
Originally inspired by the remarkable adaptability of quadruped mammal locomotion—an ability shaped by innate and environmentally induced factors [1, 2, 3]—the field of quadruped robotics has made significant efforts to develop equally proficient locomotion frameworks. Presently, the most advanced systems rely on end-to-end deep reinforcement learning (DRL), which involves training a multilayer perceptron (MLP) [4] capable of navigating diverse environments.
These frameworks demonstrate impressive realisation of terrestrial locomotion skills, which can be classified into two groups (further detailed in Supplementary Information Section 1); Froude-characterised locomotion (no desired velocity normal to the ground plane and hence upholds the assumptions of the Froude number [5]) such as walking or running in real-world [6, 7], urban [8, 9] and deformable[7] terrains, and Froude-free locomotion (features a desired velocity normal to the ground plane) such as jumping between platforms [10], climbing over obstacles [11], and sure-footedness [12]. Despite these achievements, when it comes to Froude-characterised locomotion, which can account for $\approx 70-90\%$ daily animal locomotion [13, 14], their adaptability remains constrained, as most systems are limited to deploying a single targeted gait or locomotion strategy.
Figure 1: Instilling the core animal locomotion proficiency attributes within a DRL locomotion framework. From taking an abstracted view of animal locomotion to determine the attributes of proficient locomotion, we instil animal gait transition strategies within the gait selection policy, $\pi_{G}$, gait procedural memory is embedded within the bio-inspired gait scheduler, and adaptive motion adjustments are realised by the locomotion policy, $\pi_{L}$. $\pi_{G}$ has been trained to minimse the animal gait transition metrics applied to the quadruped robot based on the current robot state, $\bm{s}$, and relevant bio-inspired gait scheduler output, $\bm{\beta}_{G}$, to select the optimal gait, $\Gamma^{*}$. The bio-inspired gait scheduler then generates the gait references informed by $\bm{s}$ from encoded high-level gait parameters. The gait references, $\bm{\beta}_{L}$, are then passed to $\pi_{L}$ to inform of any adjustments to the nominal gait motions. The control framework is constant between training and deployment to improve sim-to-real transfer.
In contrast, biomechanics research has shown that no single gait is universally optimal across all scenarios within Froude-characterised locomotion [15, 16, 17]. Animals adapt their locomotion by employing nominal gaits such as ambling, trotting, and running [18], while switching to specialised gaits like hopping, pronking, and bounding for off-nominal tasks such as predator evasion or obstacle navigation [19]. Current DRL frameworks fall short of replicating this level of Froude-characterised locomotion versatility. To address this limitation, some approaches have focused on training DRL policies to learn multiple gaits by providing reference motions during training [20, 21, 22, 23], or by learning from policies that specialise in specific gaits [24]. However, these methods remain insufficient when compared to the extensive capabilities observed in animal locomotion, which include:
@@
While bio-inspired methods that leverage central pattern generators (CPGs) [25] have realised spontaneous gait transitions [26] and mimic certain animal behaviours [27], their performance in real-world applications is often limited. Typically, such experiments are constrained to controlled environments [28, 27] or, when conducted outdoors, are restricted by low velocities and simple tasks [22]. Another bio-inspired approach is training locomotion policies from re-targeted animal motion data [29]. Although this method can achieve natural Froude-free locomotion, there is no guarantee that the exhibited Froude-characterised locomotion is optimal; it is unreasonable to assume that mechanically different systems would perform efficiently with the same low-level behaviour. These limitations suggest that instead of attempting to replicate animal locomotion mechanisms precisely, state-of-the-art DRL frameworks should be augmented with high-level attributes derived from animal locomotion to instil the proficiency observed in nature.
@@
Despite these insights, there has been no prior attempt to simultaneously integrate all these attributes—gait transition strategies, procedural memory, and adaptive motion adjustments—within existing DRL locomotion frameworks. This leads to the following research questions:
@@
To address these questions, we propose a novel DRL locomotion framework (see Figure 1) designed to incorporate these key animal locomotion attributes. The framework demonstrates exceptional adaptability through zero-shot deployment in complex, real-world environments, relying solely on interoceptive sensors.
@@ I Results
We evaluate the proposed framework through a set of studies, demonstrating it outperforms others by exhibiting quadruped animal locomotion strategies, and validating this gained proficiency on real-world terrain, as presented in Figure 2 and Supplementary Video S1.
@@ I-A Achieving Adaptive Motion Adjustment with a Diverse Set of Gaits
Figure 3: Locomotion comparison study experiments (a) each policy follows a set of command velocities in $x$ and yaw between 00 to $1$ m/s and $-1$ to $1$ rad/s respectively. During each velocity pair, the commanded gait is cycled through all gaits, switching every $1$ s. This repeated $5$ times over flat, $h_{\text{terr}}^{0}$, to very rough terrain, $h_{\text{terr}}^{3}$ (shown in Figure 2) with the average performance being plotted. A number rather than a magnitude indicates the count of experiments that the policy failed. (b) each policy follows a sinusoidal trajectory in $v_{x}$ and $\omega_{z}$ while switching gaits every $2$ s, which is repeated $5$ times over $h_{\text{terr}}^{0}$ and $h_{\text{terr}}^{3}$ terrain. Green highlighted areas indicating transition phases, for which the function and formulation are detailed in Section III-B, and black triangles representing points of failure. (c) policies follow a command of just $v_{x}=0.5$ m/s while using a constant trot gait while encountering rectangular steps with a height of $0.05$ m, where $|\bar{\omega}_{B}^{\text{err}}|$ is the magnitude of the desired orientation error.
@@
To evaluate our method of instilling adaptive motion adjustment and procedural memory for diverse gait deployment, a comparison study is completed between our bio-inspired locomotion policy $\pi_{L}^{\text{bio}}$, a standard multi-gait locomotion policy with no pseudo procedural memory $\pi_{L}^{\text{no}\bm{\beta}_{L}}$, and a policy trained which also uses $\bm{\beta}_{L}$ within $\bm{o}_{L}$ but implements the standard approach of extracting the observations from the simulator, $\pi_{L}^{\text{noSE}}$, as shown in Supplementary Video S2. From the results of this study, shown in Figure 3, the proficiency of $\pi_{L}^{\text{bio}}$ over $\pi_{L}^{\text{no}\bm{\beta}_{L}}$ and $\pi_{L}^{\text{noSE}}$ in terms of velocity tracking error $v_{B}^{\text{err}}$, contact schedule tracking error $c_{\text{avg}}^{\text{err}}$, and base stability (magnitude of undesirable base angular velocities) $\omega_{B}^{\text{err}}$, is stark.
As shown in Figure 3(a), on flat terrain, $\pi_{L}^{\text{bio}}$ achieves lower $v_{B}^{\text{err}}$, $c_{\text{avg}}^{\text{err}}$, and $\omega_{B}^{\text{err}}$ than $\pi_{L}^{\text{no}\bm{\beta}{L}}$ and $\pi{L}^{\text{noSE}}$, averaging $15\%$, $21\%$, and $10\%$ lower respectively (excluding failures), with errors increasing at higher velocity commands. This performance gap widens on rough terrain, where $\pi_{L}^{\text{no}\bm{\beta}{L}}$ and $\pi{L}^{\text{noSE}}$ fail frequently—especially at higher speeds and rougher surfaces—while $\pi_{L}^{\text{bio}}$ succeeds in all trials despite being trained only on flat terrain. This highlights its adaptability to unseen environments (see Section III-C for rationale behind omitting rough terrain in $\pi_{L}$ training).
@@
This incompetence of $\pi_{L}^{\text{no}\bm{\beta}_{L}}$ and $\pi_{L}^{\text{noSE}}$ is caused by a lack of adaptive swing foot motion adjustments captured within $\bm{\beta}_{L}$ based on $\bm{s}$ and the accumulation of error within the SE respectively. Both of these factors are substantially affected by the instabilities rough terrain inflicts upon the robot. Considering that the nominal swing foot peak height is defined as $25\%$ of the nominal base height and for $h_{\text{terr}}^{2}$ and $h_{\text{terr}}^{3}$ (as defined in Figure 2) the peak terrain height is $44\%$ and $67\%$ of the base height, having no data or strategy to account for this harsh terrain results in the rapid deterioration of proficiency.
@@
For $\pi_{L}^{\text{no}\bm{\beta}{L}}$, Figure 3(b) shows large spikes in $c_{\text{avg}}^{\text{err}}$ and $\omega_{B}^{\text{err}}$ when encountering $h_{\text{terr}}^{3}$, highlighting its inability to adapt swing foot trajectories and overcome an unrefined solution space. Figure 3(c) further shows sustained instabilities in base height and orientation after contact with steps at $17\%$ of the nominal base height. For $\pi_{L}^{\text{noSE}}$, Figs. 3(b) and 3(c) show that poor reference tracking and stability worsen with velocity command magnitude and time, as it lacks strategies to mitigate error build-up in the SE. With $\pi_{L}^{\text{bio}}$ not experiencing any of these limitations, this explicitly demonstrates the effectiveness of implementing $\bm{\beta}_{L}$ and $\bm{s}$ within $\bm{o}_{L}$; $\pi_{L}^{\text{bio}}$ successfully generalises across gaits and terrain, demonstraing successful instillation of adaptive motion adjustment and gait procedural memory.
@@ I-B Applying Biomechanics Metrics For Optimal Gait Selection
Directly applying biomechanics metrics to instil animal gait transition strategies is unsuitable due to differences between animals and robots and $\pi_{G}$ training requirements. Instead, we use cost of transport (CoT), torque saturation ($\tau_{\%}$), external work ($W_{\text{ext}}$), and foot contact tracking error ($c_{\text{avg}}^{\text{err}}$) to minimise energy use, actuator-structural forces, mechanical work, and instability, respectively. Details and justification of these metrics are in Section III-E.
@@
In accordance with Section III-F, these metrics are unified within the training of the gait selection policy $\pi_{G}^{\text{uni}}$ to instil the strategies animals use for optimal gait selection for exemplary adaptability. To investigate whether $\pi_{G}^{\text{uni}}$ effectively minimises these metrics through gait selection, the results of competing the highly demanding velocity command trajectory presented in Figure 4 and Supplementary Video S3 for $\pi_{G}^{\text{uni}}$ paired with $\pi_{L}^{\text{bio}}$ are collected, along with that for all individual gaits deployable by $\pi_{L}^{\text{bio}}$, for flat terrain and terrain $h_{\text{terr}}^{2}$.
@@
Figure 4 shows that $\pi_{G}^{\text{uni}}$ exclusively uses trotting at low speeds and running at high speeds. When accelerating, $\pi_{G}^{\text{uni}}$ oscillates between trotting and running, with an increasing bias towards running, to increase the stride frequency as visualised by the foot contact data in Figure 4. Consequently, $\pi_{G}^{\text{uni}}$ not only reliably tracks the optimal gait but outperforms individual gaits, which is only aided by its transition to standing during $v_{x}^{\text{cmd}}=0$ $\omega_{z}^{\text{cmd}}=0$ events for minimal $\tau_{\%}$, $W_{\text{ext}}$ and $c_{\text{avg}}^{\text{err}}$. This behaviour, although never targeted, is reflected in animal locomotion strategies, where gait stride frequency increases with speed, and transitional phases blend gaits to minimise energy use [33]. Under rough terrain and high acceleration, $\pi_{G}^{\text{uni}}$ employs additional gaits to manage instability. This leads to a gait classification: trot and run serve as nominal gaits for low and high speeds, while bound, pronk, limp, amble, and hop act as auxiliary gaits for off-nominal conditions like stability recovery.
@@
When inspecting the radar charts in Figure 4, which depict relative performance in terms of the metrics, the origin of this emerged gait selection strategy becomes clear. Across the gaits, on flat terrain trot and run gaits achieve the best relative performance. However, when it comes to overcoming rough terrain bound, hop and limp gaits all gain relative performance in $\tau_{\%}$, $W_{\text{ext}}$ and $c_{\text{avg}}^{\text{err}}$, while trot and run gaits exhibit a reduced dominance in relative proficiency. In addition to this observation providing an insight as to why $\pi_{G}^{\text{uni}}$ chooses to utilise these auxiliary gaits it also provides evidence to suggest that $\tau_{\%}$, $W_{\text{ext}}$ and $c_{\text{avg}}^{\text{err}}$ can effectively characterise stability. This observation is further investigated and discussed in the following sections. Overall, $\pi_{G}^{\text{uni}}$ outperforms all individual gaits across all metrics, with the exception of CoT for trot and run gaits where the difference is negligible, demonstrating the successful minimisation of the metrics and successful instillation of gait procedural memory of how to utilise each gait given the robot’s state and task.